OpenStudy (anonymous):

Solve the polynomial function below. -192x^2 + 27 = 0

OpenStudy (solomonzelman):

1) subtract 27 from both sides \[-192x^2 = -27\] 2) divide both sides by -3 \[64x^2 = 9\] 3) divide both sides by 64 \[x^2 = \frac{9}{64}\] 4) square-root both sides \[x =±\sqrt{ \frac{9}{64} }\] can you finish the problem ?

OpenStudy (anonymous):

Ive never really done this before so its a bit hard...

OpenStudy (solomonzelman):